OpenStudy (anonymous):

Use completing the square to solve: 4x^2+4x-3=0

OpenStudy (anonymous):

\[4x ^{2}+4x-3=0\] \[x ^{2}+x-3/4=0\] \[x ^{2}+2x(1/2)+(1/2)^{2}-(1/2)^{2}=3/4\] \[(x+1/2)^{2}=3/4+1/4\] \[(x+1/2)^{2}=1\] \[x+1/2=1 or x+1/2=-1\] \[\implies x=1/2,-3/2\] are the 2 values of x

OpenStudy (anonymous):

so half of x is 2x?

OpenStudy (anonymous):

\[4(x^2+x-\frac{3}{4})=0\]\[4[(x+\frac{1}{2})^2-\frac{1}{4}-\frac{3}{4}]=0\]\[(x+\frac{1}{2})^2-1=0\]\[(x+\frac{1}{2})^2=1\]\[x=\pm 1-\frac{1}{2}\]

OpenStudy (anonymous):

@sarah you got excatly the same answer as i got

OpenStudy (anonymous):

@chrissy who said that?

OpenStudy (anonymous):

@luckey yeah i know i used a different method

OpenStudy (anonymous):

dont you take half of the x term then square it??

OpenStudy (anonymous):

@sarah both the methods takes the same time right........ i went through your answer

OpenStudy (anonymous):

right:)

OpenStudy (anonymous):

@chrissy thats not taking half of x and squaring it.... that was just an arrangement to make it a perfect square as; \[(a+b)^{2}=a ^{2}+2ab+b ^{2}\] got it??

OpenStudy (anonymous):

sweet, thanks guys and dolls.

OpenStudy (anonymous):

@chrissy you're always welcomed :))))

OpenStudy (anonymous):

actually how did you get that 2x?

OpenStudy (anonymous):

thats not 2x., thats still x only \[2x(1/2)=x \]

OpenStudy (anonymous):

so x(1/2)= 1/2

OpenStudy (anonymous):

because the x is just 1

OpenStudy (anonymous):

no no thats x(1/2) only ok look 2x(1/2)=2(1/2)x 2 2 cancelled then we r left with just x

OpenStudy (anonymous):

you said before that 2x is still only x, how is that possible?

OpenStudy (anonymous):

you please go through the solution once again properly

OpenStudy (anonymous):

@sarah please come and explain her yaar

OpenStudy (anonymous):

if I have £2 i have two £1 right?

OpenStudy (anonymous):

so how can 2x remain x??? if I am being told that there are 2x's

OpenStudy (anonymous):

the other ladies working show no 2x

OpenStudy (anonymous):

\[x^2+x=\frac{3}{4}\] and you are now thinking that the coefficient of the "x" term is 1, as in \[x^2+1x=\frac{3}{4}\]

OpenStudy (anonymous):

so you think the following: 1) half of 1 is 1/2 2) one half squared is 1/4

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and then write \[(x+\frac{1}{2})^2=\frac{3}{4}+\frac{1}{4}=1\]

OpenStudy (anonymous):

then you write \[x+\frac{1}{2}= 1\] or \[x+\frac{1}{2}=-1\]

OpenStudy (anonymous):

and finally \[x=1-\frac{1}{2}=\frac{1}{2}\] OR \[x=-1-\frac{1}{2}=-\frac{3}{2}\]

OpenStudy (anonymous):

if you want we can do another example

OpenStudy (anonymous):

so what about the 2x that is not really 2x but just x business??

OpenStudy (anonymous):

yes please another example would be great.

OpenStudy (anonymous):

ok i do one of my choosing and write down exactly what i am thinking. start with an easy one \[x^2+6x-1=0\]

OpenStudy (anonymous):

add one to both sides \[x^2+6x=1\]

OpenStudy (anonymous):

would you mind if I tried to solve it first.

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

I am stuck. i got to (x+3)=10

OpenStudy (anonymous):

ok it goes like this \[(x+3)^2=10\] i think you forgot the square, but the rest is right

OpenStudy (anonymous):

then i write \[x+3=\sqrt{10}\] or \[x+3=-\sqrt{10}\] and finally \[x=-3+\sqrt{10}\] or \[x=-3-\sqrt{10}\]

OpenStudy (anonymous):

how about \[x^2+4x-2=0\]

OpenStudy (anonymous):

thats is where I got stuck, 10 is not a square number so my brain froze.

OpenStudy (anonymous):

ah ok. well we just pretend like we know what were are doing and write \[\sqrt{10}\] if you get a perfect square, like say 4 or 9 or 16, then you know what to write

OpenStudy (anonymous):

you can practice with \[x^2+2x-3=0\] if you want. that will give you a perfect square

OpenStudy (anonymous):

(x+2)^2=6 (x+2)=6???

OpenStudy (anonymous):

you got it, but make sure if you get \[(x+2)^2=6\] to write \[x+2=\sqrt{6}\text { or } x+2=-\sqrt{6}\]

OpenStudy (anonymous):

x=-3 and 1

OpenStudy (anonymous):

for the second question

OpenStudy (anonymous):

for the second one yes. good work! did you "complete the square"?

OpenStudy (anonymous):

(x-1)^2

OpenStudy (anonymous):

sorry (x+1)^2

OpenStudy (anonymous):

got it. here is a harder one, but still perfect square answer \[2x^2+x-1=0\] this one is tricky so be careful

OpenStudy (anonymous):

you where right this one is trick, the lone x term as really thrown me off.

OpenStudy (anonymous):

still trying thought. gimme another 10 mins lol

OpenStudy (anonymous):

yeah and what is even trickier is that you have to divide by twice! once to get the "leading coefficient " to be 1 and again when you complete the square

OpenStudy (anonymous):

i got x = -9/4 and 7/4 :( and the square (x+1/4)^2

OpenStudy (anonymous):

the square part is right. so you got one one side \[(x+\frac{1}{4})^2\] and that is correct. what did you get on the other side?

OpenStudy (anonymous):

8/16

OpenStudy (anonymous):

ah you are off by one

OpenStudy (anonymous):

it should be \[x^2+\frac{1}{2}x=\frac{1}{2}\] then \[(x+\frac{1}{4})^2=\frac{1}{2}+\frac{1}{16}\]

OpenStudy (anonymous):

and \[\frac{1}{16}+\frac{1}{2}=\frac{9}{16}\]

OpenStudy (anonymous):

now when you take the square root, don't forget to take the square root of numerator and denominator, so you should have \[x+\frac{1}{4}=\frac{3}{4}\] or \[x+\frac{1}{4}=-\frac{3}{4}\]

OpenStudy (anonymous):

what If i get a situation where I the numerator is a root and the denominator is not

OpenStudy (anonymous):

like 8/9

OpenStudy (anonymous):

you mean the other way around. no problem \[\sqrt{\frac{8}{9}}=\frac{\sqrt{8}}{\sqrt{9}}=\frac{\sqrt{8}}{3}\]

OpenStudy (anonymous):

that is what you usually get, because the denominator you get from squaring, so it is usually a perfect square to begin with. it is the numerator that is often not a square

OpenStudy (anonymous):

thank you for all you help satellite

OpenStudy (anonymous):

yw